OpenStudy (anonymous):

simplify (29/6)+(17/10)+x+(37/8)

OpenStudy (anonymous):

\[\frac{29}{6}+ \frac{17}{10}+x+\frac{37}{8}\] Next we will try to make the denominators of each fraction equal.

OpenStudy (anonymous):

\[=\frac{29 \times 8 \times 10}{6\times 8 \times 10}+ \frac{17\times 8 \times 6}{10\times 8 \times 6}+x+\frac{37\times 6 \times 10}{8\times 6 \times 10}\] \[=\frac{2320}{480}+ \frac{816}{480}+x+\frac{2220}{480}=\frac{2320+816+2220}{480}+x\] \[=\frac{5356}{480}+x= 11.1583+x\]

OpenStudy (anonymous):

ok cool thanks but there are plenty more where that came from